# MdEval: Massively Multilingual Code Debugging

### v1

@@ 5 Related Work
With the rapid advancement of large language models(LLMs) (OpenAI, 2023; Touvron et al., 2023b; AI, 2024; Bai et al., 2023; Yang et al., 2024a), solving complex code-related tasks has become increasingly feasible, leading to the emergence of numerous Code LLMs. Early studies utilized models like BERT (Devlin et al., 2019) or GPT (Radford et al., 2018) as backbones, trained on billions of code snippets to enable tasks involving code understanding and generation (Chen et al., 2021; Feng et al., 2020; Scao et al., 2022; Li et al., 2022; Wang et al., 2021; Allal et al., 2023). Recently, advancements in domain-specific pre-training and instruction fine-tuning techniques (Zheng et al., 2024a; Yue et al., 2024) have led to extensive efforts in fine-tuning models on large-scale code corpora and crafting code-related task instructions (Rozière et al., 2023; Zheng et al., 2023; Luo et al., 2023; Muennighoff et al., 2023; Gemma Team, 2024; Zheng et al., 2024b; Guo et al., 2024a; Wei et al., 2023; Sun et al., 2024; Lozhkov et al., 2024; Jiang et al., 2023; Hui et al., 2024b; Wang et al., 2024a; Deng et al., 2024; Liu et al., 2024). These models demonstrate remarkable performance in tasks like code completion, synthesis, and program repair.
@@
Automatic program debugging holds substantial practical value. With the emergence of LLM capabilities, a growing number of individuals are utilizing LLMs for code debugging, leading to extensive research in this field. Code Debugging includes serval tasks such as bug or vulnerability detection (Pradel and Sen, 2018; Allamanis et al., 2021; Yuan et al., 2023; Zhang et al., 2024; Zhong et al., 2024), fuzz test (Deng et al., 2023; Xia et al., 2024; Yang et al., 2024b), program repair  (Wen et al., 2024; Lin et al., 2017; Zhang et al., 2023a; Prenner and Robbes, 2023; Gu et al., 2024; Tambon et al., 2024; Wang et al., 2024b), GitHub issues auto resolving  (Jimenez et al., 2023; Chen et al., 2024; Tao et al., 2024). To effectively assess the code debugging capabilities of LLMs, several benchmark tests have been introduced  (Prenner et al., 2022; Sobania et al., 2023; Xia and Zhang, 2023; Zhang et al., 2023b; Tian et al., 2024; Yang et al., 2024c). Notably, DebugBench (Tian et al., 2024) provides a comprehensive classification of error types and analyzes the debugging capabilities of LLMs based on these categories. Similarly, DebugEval (Yang et al., 2024c) has designed various debugging-related tasks to evaluate LLM performance across different task dimensions. However, these studies focus on 1 to 3 languages. In reality, there are significant differences in code errors between languages, leading to numerous language-specific errors. To address this gap, we propose MdEval, a comprehensive code debugging benchmark covering 18 languages, aiming to assess LLM debugging capabilities from a broader perspective.
@@ 6 Conclusion
In this work, we introduce MdEval of instruction corpora MdEval-Instruct, evaluation benchmark, and a strong baseline xDebugCoder, where the benchmark includes automated program repair (APR), code review (CR), and bug identification (BI) of 18 programming languages (total 3.6K samples), aiming to assess the debugging capabilities of large language models (LLMs) in multilingual environments. Further, we propose xDebugGen to construct a multilingual debugging instruction corpus, where we inject the bugs into the query or answer to create the pair of the buggy code and correct code. Based on MdEval-Instruct, we develop xDebugCoder, a multilingual LLM for debugging in a wide range of programming languages as a strong baseline. Through extensive experiments, this paper reveals a substantial performance gap between open-source and closed-source LLMs, underscoring the need for further improvements in multilingual code debugging. In the future, we will continue expanding the number of languages in MdEval.

### v2

@@ 5 Related Work
The rapid progress of large language models(OpenAI, 2023; Touvron et al., 2023b; AI, 2024; Bai et al., 2023; Yang et al., 2024a) has enabled complex code-related tasks. Early models like BERT(Devlin et al., 2019) and GPT (Radford et al., 2018), trained on billions of code snippets, focused on code understanding and generation (Chen et al., 2021; Feng et al., 2020; Scao et al., 2022; Li et al., 2022; Wang et al., 2021; Allal et al., 2023). Recent advances in domain-specific pre-training and instruction fine-tuning (Zheng et al., 2024a; Yue et al., 2024) have enhanced models like CodeLlama (Rozière et al., 2023) and WizardCoder (Luo et al., 2023), achieving strong performance in code completion, synthesis, and repair.
@@
LLMs have also gained popularity for automatic program debugging, a critical task for bug detection, vulnerability identification (Pradel and Sen, 2018; Allamanis et al., 2021), fuzz testing (Deng et al., 2023; Xia et al., 2024), and program repair (Wen et al., 2024; Gu et al., 2024). Benchmark tests, such as DebugBench (Tian et al., 2024) and DebugEval (Yang et al., 2024c), assess LLM debugging capabilities across error categories and tasks. However, these focus on 1-3 languages, neglecting language-specific errors. To fill this gap, we propose MdEval, a comprehensive debugging benchmark for 20 languages to evaluate LLM performance from a broader perspective.
@@ 6 Conclusion
In this work, we introduce MdEval of instruction corpora MdEval-Instruct, evaluation benchmark, and a strong baseline xDebugCoder, where the benchmark includes automated program repair (APR), bug localization (BL), and bug identification (BI) of 20 programming languages (total 3.9K samples), aiming to assess the debugging capabilities of large language models (LLMs) in multilingual environments. Further, we propose xDebugGen to construct a multilingual debugging instruction corpus, where we inject the bugs into the query or answer to create the pair of the buggy code and correct code. Based on MdEval-Instruct, we develop xDebugCoder, a multilingual LLM for debugging in a wide range of programming languages as a strong baseline. Through extensive experiments, this paper reveals a substantial performance gap between open-source and closed-source LLMs, underscoring the need for further improvements in multilingual code debugging. In the future, we will continue expanding the number of languages in MdEval.
Appendix A Limitations
Language Coverage.
Although MdEval covers 20 programming languages, there are still many languages not included, particularly those that are less commonly used or have niche applications. Expanding the benchmark to include more languages would provide a more comprehensive evaluation of multilingual debugging capabilities.
Real-world Applicability.
While MdEval aims to simulate realistic debugging scenarios, the tasks and data may not fully capture the complexity and variability of real-world software development. Incorporating more diverse and complex real-world projects into the benchmark could improve its applicability and relevance.
Instruction Tuning Data.
The instruction corpora MdEval-Instruct used for fine-tuning the baseline model xDebugCoder is generated by LLM-based bug injection. While this approach has shown promise, the quality and diversity of the generated data could be further improved. Exploring alternative methods for generating high-quality instruction data, such as leveraging more advanced LLMs or incorporating feedback from real-world debugging sessions, could enhance the effectiveness of the instruction tuning process.
Appendix B Ethical Considerations
B.1 Potential Risks
MdEval, as evaluation tools, can comprehensively assess the capability of large language models in debugging tasks across a wide range of programming languages, thereby advancing the development of large language models in this domain. However, improper or erroneous use of MdEval may pose significant risks, such as incorrect program analysis and faulty program repair, which could even lead to severe consequences such as program crashes or operating system failures. Therefore, to ensure the security and reliability of the evaluation process, we strongly recommend using MdEval within a sandbox environment. Such an environment can effectively isolate potential system risks, ensuring the accuracy and safety of the evaluation.
Appendix C Human Annotation
To construct the massively multilingual code debugging benchmark MdEval, we designed and implemented a comprehensive and systematic human annotation process to ensure the accuracy, consistency, and high quality of multilingual code samples. This process strictly adheres to carefully formulated annotation guidelines and incorporates multiple quality control mechanisms.
We recruited 13 computer science graduates as multilingual debugging annotators, all of whom are proficient in at least one programming language and possess a solid foundation in computer science. Prior to the formal annotation process, annotators underwent systematic training on annotation methods, covering core tasks such as problem definition, solution design, and buggy code generation.
Our annotation training guidelines focus on the following key aspects:
  • Standardized Format: We provide detailed annotation examples and templates for 20 programming languages. Annotators must strictly adhere to a standardized format throughout the annotation process to ensure data consistency and reusability.
  • Accessibility: All annotation reference data are sourced from open-source materials that allow free use and distribution, ensuring compliance with academic research purposes and relevant legal and ethical requirements.
  • Difficulty Classification: We establish a detailed difficulty classification guideline for each programming language. Annotators must categorize each problem according to complexity, error type, and problem scale, assigning an appropriate difficulty level (e.g., easy, middle, hard) following the guidelines.
  • Self-Containment: Annotators must ensure that each problem description is complete and unambiguous, containing all necessary information for problem-solving. Provided example inputs and outputs must be accurate, the generated buggy code must be ensured to fail execution correctly, and the reference solution must pass all test cases. Additionally, test cases should comprehensively cover various boundary conditions and exceptional scenarios.
To maintain annotation quality and incentivize annotators, we offered a compensation of approximately \$6 per problem. Moreover, we provided annotators with a comfortable working environment, free meals, souvenirs, and high-performance computing equipment. A total of approximately 1,300 problems were annotated, with additional annotators hired for quality inspection, leading to a total cost of around \$5,000. Quality inspection tasks included bug identification, bug localization, and code review.
C.1 Quality Control
To ensure the high quality of the MdEval, we implemented a rigorous quality control mechanism. First, annotators were required to evaluate the annotated code based on four core criteria: problem difficulty, ambiguity, error type, and solvability. Second, we adopted a dual verification system, where each code snippet was independently annotated by at least two annotators to minimize subjective bias and human errors. In cases of disagreement, resolution was achieved through discussion or by a senior annotator making the final decision.
To further ensure the reliability of the benchmark, we employed three volunteers to assess whether MdEval achieved a correctness rate of at least 90% and to correct any errors, thereby guaranteeing the accuracy of the annotations.
Appendix D Experiment Detail
xDebugCoder Training Corpora.
xDebugCoder Optimization.
Our model, xDebugCoder, based on Qwen2.5-Coder-7B, is trained for 3 epochs using a cosine scheduler, starting at a learning rate of $5\times 10^{-5}$ with 3% of total training steps for warmup. We utilize AdamW (Loshchilov and Hutter, 2017) as the optimizer; the batch size is set to 1024, with a maximum sequence length of 2048. All experiments are performed with 8 NVIDIA A800-80GB GPUs.
Code LLMs.
We evaluate 40 popular models, both closed-source and open-source (sizes ranging from 1.3B to 605B parameters). For general models, we evaluate GPTs (OpenAI, 2023) (GPT4-o, GPT4-o-mini), Claude-3.5 (Anthropic, 2023). For code models, we test Qwen2.5-Coder (Hui et al., 2024), DeepSeekCoder (DS-Coder) (Guo et al., 2024a), CodeLlama (Rozière et al., 2023), and Codegemma (Gemma Team, 2024). Furthermore, we fine-tune the Qwen2.5-Coder-7B to provide a baseline model xDebugCoder for reference. For closed-source models, the responses are generated by the official API. For the open-source models, we perform inference on all models using the vLLM (Kwon et al., 2023) framework. All models adopt a greedy decoding strategy during inference, the temperature is set to 0, and the maximum generation length is 4096.
Appendix E Related Work
Code Large Language Model.
With the rapid advancement of large language models(LLMs) (OpenAI, 2023; Touvron et al., 2023b; AI, 2024; Bai et al., 2023; Yang et al., 2024a), solving complex code-related tasks has become increasingly feasible, leading to the emergence of numerous Code LLMs. Early studies utilized models like BERT (Devlin et al., 2019) or GPT (Radford et al., 2018) as backbones, trained on billions of code snippets to enable tasks involving code understanding and generation (Chen et al., 2021; Feng et al., 2020; Scao et al., 2022; Li et al., 2022; Wang et al., 2021; Allal et al., 2023). Recently, advancements in domain-specific pre-training and instruction fine-tuning techniques (Zheng et al., 2024a; Yue et al., 2024) have led to extensive efforts in fine-tuning models on large-scale code corpora and crafting code-related task instructions (Rozière et al., 2023; Zheng et al., 2023; Luo et al., 2023; Muennighoff et al., 2023; Gemma Team, 2024; Zheng et al., 2024b; Guo et al., 2024a; Wei et al., 2023; Sun et al., 2024; Lozhkov et al., 2024; Jiang et al., 2023; Hui et al., 2024; Wang et al., 2024a; Deng et al., 2024; Liu et al., 2024). These models demonstrate remarkable performance in tasks like code completion, synthesis, and program repair.
Debugging with Large Language Models.
Automatic program debugging holds substantial practical value. With the emergence of LLM capabilities, a growing number of individuals are utilizing LLMs for code debugging, leading to extensive research in this field. Code Debugging includes serval tasks such as bug or vulnerability detection (Pradel and Sen, 2018; Allamanis et al., 2021; Yuan et al., 2023; Zhang et al., 2024; Zhong et al., 2024), fuzz test (Deng et al., 2023; Xia et al., 2024; Yang et al., 2024b), program repair  (Wen et al., 2024; Lin et al., 2017; Zhang et al., 2023; Prenner and Robbes, 2023; Gu et al., 2024; Tambon et al., 2024; Wang et al., 2024b), GitHub issues auto resolving  (Jimenez et al., 2023; Chen et al., 2024; Tao et al., 2024). To effectively assess the code debugging capabilities of LLMs, several benchmark tests have been introduced  (Prenner et al., 2022; Sobania et al., 2023; Xia and Zhang, 2023; Zhang et al., 2023; Tian et al., 2024; Yang et al., 2024c). Notably, DebugBench (Tian et al., 2024) provides a comprehensive classification of error types and analyzes the debugging capabilities of LLMs based on these categories. Similarly, DebugEval (Yang et al., 2024c) has designed various debugging-related tasks to evaluate LLM performance across different task dimensions. However, these studies focus on 1 to 3 languages. In reality, there are significant differences in code errors between languages, leading to numerous language-specific errors. To address this gap, we propose MdEval, a comprehensive code debugging benchmark covering 20 languages, aiming to assess LLM debugging capabilities from a broader perspective.